# CT-UIO: Continuous-Time UWB-Inertial-Odometer Localization Using Non-Uniform B-spline with Fewer Anchors

### v1

@@ III System Overview
Fig. 1 shows the overviews of our CT-UIO. In the beginning, we assume that the UWB-Inertial-Odometer system is calibrated and IMU bias and odometer scale are initialized. In the preprocessing, we apply inequality constraints to remove UWB-ranging outliers and build the IMU/odometer fusion model (Sec. III-A). Next, we propose a non-uniform continuous-time state estimation method, using the adaptive knot span adjustment strategy to tightly fuse the high-frequency measurements from UWB ranging, IMU, and odometer (Sec. III-B). To make the few-anchor UWB localization system fully observable, we combine the UWB ranging measurements with short-term position estimates from the IMU/odometer fusion model to generate virtual UWB anchors (Sec. III-C). In the backend, we conduct a UWB-Inertial-Odometer factor graph with an adaptive sliding window to perform joint optimization to obtain optimal estimations (Sec. III-D). For UWB/IMU/Odometer integrated (UIO) system, we use ${U}$ to denote the UWB frame, $I$ for the IMU frame, $O$ for the wheel odometer frame, $W$ for the world coordinate frame.
III-A Preprocessing
III-A1 IMU/Odometer Fusion Model
@@
IMU and odometer offer fast response speeds and are environment-independent. For ground robot navigation, it appears to be straightforward in to fuse IMU data with wheel odometer measurements to provide motion priors over short time periods [30]. The raw IMU measurements include the angular velocity $\omega_{t}$ and local linear acceleration $a_{t}$
$\omega_{t_{k}}=\tilde{\omega}_{t_{k}}+\varepsilon_{g}+n_{g},$ (1)
$a_{t_{k}}=\tilde{a}_{t_{k}}+\varepsilon_{a}+\prescript{I}{W}{\textbf{R}}g^{W}+n_{a},$ (2)
@@
where $\tilde{\omega}_{t_{k}}$ and $\tilde{a}_{t_{k}}$ are true values of angular velocity and linear acceleration at time $t$, respectively. $\varepsilon_{g}$ and $\varepsilon_{a}$ denote accelerometer’s bias and gyroscope’s bias in the IMU body frame. $\prescript{I}{W}{\textbf{R}}$ denotes the rotation matrix from frame $\left\{W\right\}$ to $\left\{I\right\}$, $g^{W}$ is the gravity vector in the world frame, $n_{g}$ and $n_{a}$ are zero-mean Gaussian noises. We then integrate the wheel odometer measurements between IMU times [$t_{k-1}$, $t_{k}$] and 2D relative motion as follows:
[…]
@@
$\tilde{z}_{o}=:\begin{bmatrix}{}_{I}^{o}\theta\\ {}_{I}^{o}\textbf{d}\par\end{bmatrix}=\begin{bmatrix}\textbf{e}_{3}^{\text{T}}{{}^{o}R_{t_{k}}}\\ {{}^{o}R_{t_{k}}}(_{I}^{o}R{{}^{I}\textbf{d}})\end{bmatrix}.$ (4)
The attitude update is formulated as
@@
${{}^{o}R_{t_{k}}}$ represents the rotation matrices transformed from ${}^{o}\theta_{t_{k}}$, $\textbf{e}_{3}$ is the unit bias vector. We construct the observation measurement vector composed of the pose of IMU and odometer as
@@
where $\eta(k)\sim(0,\sigma^{2})$ means the measurement noise with zero mean and standard Gaussian distribution. The Jacobian matrix $H(k)$ can be indicated by
@@
The innovation $\tilde{Z}_{k}$ can be calculated as
@@
where $M$ is the adaptive size of the observation statistics, which controls the trade-off between estimation accuracy and computational complexity. $\mu$ is the convergence rate. There is the rejection domain of the $E(k)$
$\left\{\begin{matrix}E(k)>thr&\text{Reject}\\ E(k)\leq thr&\text{Accept}\end{matrix}\right.$ (14)
@@
when $E(k)$ exceeds the threshold $thr$, the abnormal fault in IMU measurements is detected by chi-square test statistic, we reject this IMU measurement and do not perform the EKF update.
III-A2 UWB Ranging Outlier Rejection
$\sigma_{r}<\left|{\left\|\hat{d}_{t_{k}}-\textbf{p}_{n}\right\|}_{2}-r_{t_{k+1}}\right|<\gamma\hat{v}_{t_{k}}\left(t_{k+1}-t_{k}\right),$ (15)
where $\hat{d}_{t_{k}}$ and $\hat{v}_{t_{k}}$ are the estimated robot’s position and velocity derived from the IMU/odometer fusion model, respectively. $\textbf{p}_{n}$ is the position of the fixed anchor. $\sigma_{r}$ denotes the confidence intervals for the individual ranging measurements, $\gamma$ is the outlier rejection gating threshold.
III-B Adaptive Knot Span Adjustment Strategy
III-B1 Non-Uniform Continuous-Time Trajectory Representation
@@
In this paper, we use the split non-uniform B-spline representation to separately represent the translation part $\textbf{p}(t)$ in $\mathbb{R}^{3}$ and the rotation part $\textbf{R}(t)$ in the Lie group $SO(3)$ with two splines. The sampled translation vector in trajectory of the order $k$ B-spline at time $t\in[t_{k},t_{k-1}]$, can be denoted as:
$\textbf{p}(t)=\textbf{p}_{i}+\sum_{j=1}^{k}\lambda_{j}(t)\cdot\textbf{d}_{j}^{i},$ (16)
@@
where $\textbf{p}_{i}$ denotes the $i$-th control point, $\textbf{p}_{i}\in\mathbb{R}^{3}$, $t\in T=\left\{t_{0},t_{1},...,t_{n}\right\}$ is the arbitrarily independent knot time of B-splines. The coefficients ${\lambda}_{j}(t)$ represents the $j$th column of $\boldsymbol{\lambda}(t)$. $\tilde{\boldsymbol{\Lambda}}^{(k+1)}(i)$ is a non-constant cumulative basis matrix which is defined by its order $k$. $\textbf{u}(t)=\begin{bmatrix}1&u(t)&...&u^{k}(t)\end{bmatrix}^{\text{T}}$ is the normalized time vector:
@@
$\textbf{d}_{j}^{i}=\textbf{p}_{i+j}-\textbf{p}_{i+j-1}.$ (19)
@@
To balance the overall accuracy and convergence speed, we exploit a cumulative cubic B-spline function ($k=3$) in this article, the order of the spline is set to $4$.
III-B2 Adaptive Knot Spans Adjustment
For uniform-spline-based methods, control points are temporally uniformly distributed, knot spans between them are equal to construct the B-spline. However, this uniform placement does not guarantee optimal trajectory interpolation performance, especially when encoding time-varying states. We argue that it is reasonable to reduce redundant control points for steady movement. In addition, when the robot’s motion with significant velocity fluctuations over short time intervals, a finer granularity of knots is required to capture rapid changes, which requires inserting additional control points. To solve this problem, we propose an adaptive knot span adjustment strategy that dynamically adjusts the distribution of control points to flexibly accommodate varying motion speeds, as shown in Fig. 2. Namely, the current pose is registered as a Key Pose (KP) if the tracked trajectory segment has a local linear velocity or angular velocity change above a certain threshold. Specifically, we obtain the linear velocity and angular velocity respectively in a sliding window:
$v(T_{i-1},T_{i})$ $=\frac{1}{n(T_{i-1},T_{i})}\sum_{k\in n(T_{i-1},T_{i})}$ (21)
@@ III-B2 Adaptive Knot Spans Adjustment
We associate each KP with the number of control points $n_{cp}$
$n_{cp}=\begin{cases}1,&\text{if }v(T_{i-1},T_{i})\in[0,v_{\max}]\\ &\text{or }\omega\in[0,\lambda_{1}\omega_{\max}],\\ 2,&\text{if }v\in[\theta_{1}v_{\max},\theta_{2}v_{\max}]\\ &\text{or }\omega\in[\lambda_{1}\omega_{\max},\lambda_{2}\omega_{\max}],\\ 3,&\text{if }v\in[\theta_{2}v_{\max},\theta_{3}v_{\max}]\\ &\text{or }\omega\in[\lambda_{2}\omega_{\max},\lambda_{3}\omega_{\max}],\\ 4,&\text{if }v\in[\theta_{3}v_{\max},v_{\max}]\\ &\text{or }\omega\in[\lambda_{3}\omega_{\max},\omega_{\max}].\end{cases}$ (23)
We know that the knot span is actually influenced by the number of control points at a specific time interval. Each of the knot spans is independent of others, and knot spans are divided by control points
$\bigtriangleup t_{k}=\frac{1}{n_{cp}}(T_{i}-T_{i-1}).$ (24)
@@
Figure 2: Illustrations of the proposed adaptive knot span adjustment strategy. The top shows the robot’s velocity and the bottom shows that the distribution of control points is adaptive to speed variations.
Then, the trajectory segment over the sliding window $[T_{i-1},T_{i}]$ consists of $n_{cp}$ piecewise functions. We can query the pose with arbitrary timestamp $t_{i}$ in this segment. The control points of the active trajectory segment are further optimized and updated within the sliding window by a factor graph optimization. Meanwhile, control points outside the active segment remain static in the optimization but still retain marginalization priors for the next sliding window.
III-C UWB-Inertial-Odometer Factor-Graph Framework
III-C1 Virtual UWB Anchor Ranging Factor
Since directly fusing range measurements can lead to unobservability when only one or two anchors are available [2], we utilize Virtual Anchors (VA) to construct a fully observable UWB-based localization system. Inspired by [32, 33], we propose a VA generation method based on multiple hypotheses. This method introduces location reference points that are virtually placed in the UWB network, providing additional constraints for the robot’s global localization, as shown in Fig. 3. The VA is established by a location estimator from the IMU/odometer fusion model. In this paper, considering a mobile robot equipped with a UWB tag that receives a signal from a UWB anchor, the corresponding virtual anchor can be generated by range measurements and the estimated position using the IMU/odometer fusion model. The range residual is established as
$\textbf{e}_{j}^{r}=r_{j}-\left\|\textbf{x}_{j}-\textbf{p}_{n}\right\|_{2},$ (25)
Figure 3: Illustrations of the Virtual Anchor (VA) generation process.
Since $\textbf{q}^{\omega}$ and $\triangle\textbf{d}_{ij}$ can be derived from Eq. (3), the robot’s pose and derived the relative position can be expressed as the virtual anchor’s position. Eq. (25) can be further reformulated as follows:
where $\hat{\textbf{p}}_{n}^{j}$ is regarded as the position of the virtual UWB anchor. Based on this, the virtual anchor is generated through the ranging-gathering process that involves a set of robot waypoints. However, the random waypoints are often ill-posed in practice, resulting in suboptimal solutions. Selecting a proper information metric and determining the optimal waypoint placement is critical for improving accuracy. The Fisher information Matrix (FIM) [34] is considered to evaluate the performance of VA’s position estimation, and it is defined as:
$\textbf{F}_{i}=\frac{1}{\sigma_{r}^{2}}\textbf{J}^{\text{T}}\textbf{J},$ (28)
where $\sigma_{r}$ is the standard deviation of the ranging measurements. J represents the Jacobian matrix of the ranging measurement model with respect to anchor’s position $\hat{\textbf{p}}_{n}^{j}$. The determinant of the FIM $det(\textbf{F}_{i})$ contains all the ranging measurement received during time $i$ to $j$. We can set a threshold $\tau$ and generate the virtual UWB anchor when
$det(\textbf{F}_{i})>\tau.$ (30)
By maximizing $det(\textbf{F}_{i})$, the lower bound of the VA’s position estimation error can be directly determined and the optimal VA placement can be decided.
III-C2 Multiple Hypotheses-based Criteria for Virtual Anchors
We present the multiple hypotheses-based criteria to avoid collineation of generated virtual anchors. The virtual anchor is derived from the IMU/odometer fusion model, if all the VA’s 2D positions constitute a straight line, it becomes difficult to obtain a unique solution for the target’s position without additional constraints. To mitigate this problem, we use the determinant to ensure that generated virtual anchors are non-collinear. Assuming the locations of three virtual UWB anchors are $\hat{\textbf{p}}_{n}^{j}=(\hat{x}_{n}^{j},\hat{y}_{n}^{j}),\hat{\textbf{p}}_{n}^{j1}=(\hat{x}_{n}^{j1},\hat{y}_{n}^{j1}),\hat{\textbf{p}}_{n}^{j2}=(\hat{x}_{n}^{j2},\hat{y}_{n}^{j2})$. The determinant $D$ is defined as:
$D=\begin{vmatrix}\hat{x}_{n}^{j}&\hat{y}_{n}^{j}&1\\ \hat{x}_{n}^{j1}&\hat{y}_{n}^{j1}&1\\ \hat{x}_{n}^{j2}&\hat{y}_{n}^{j2}&1\end{vmatrix}.$ (31)
If $D=0$, the three virtual UWB anchors are collinear. We use a sample circle with multiple virtual anchor position hypotheses to avoid collinearity. Thus, assuming $h_{j}$ different hypotheses for position $\hat{\textbf{p}}_{n}^{j}$, the UWB virtual anchor position will be written in the following form:
$\textbf{b}_{n}^{j}=\textbf{H}_{j}^{k}\hat{\textbf{p}}_{n}^{j}$ (32)
$\textbf{H}_{j}^{k}=\begin{bmatrix}r_{j}&cos(2\pi(k-1)/h_{j})&0\\ 0&r_{j}sin(2\pi(k-1))/h_{j})&0\\ 0&0&1\end{bmatrix},k=1,...,h_{j}$ (33)
the resultant virtual anchor positions are obtained by minimizing the following expression:
arg $\text{arg min}\sum_{j\in\Omega}\frac{1}{h_{j}}\sum_{m=1}^{h_{j}}(r_{j}-\left\|\textbf{x}_{i}-\textbf{b}_{n}^{jm}\right\|)$ (34)
where $\Omega$ is the set of location information of the generated VAs.
III-C3 IMU Factor

### v2

@@ IV System Overview
Fig. 1 shows the overview of our CT-UIO. In the beginning, we assume that the UWB-Inertial-Odometer system is calibrated, and the IMU bias and odometer scale are initialized. In the preprocessing, we apply inequality constraints to remove UWB-ranging outliers and build the IMU/odometer fusion model (Sec. IV-A). To make the few-anchor UWB localization system fully observable, we combine the UWB ranging measurements with short-term position estimates from the IMU/odometer fusion model to generate VAs (Sec. IV-B1). Furthermore, we present multiple hypotheses-based criteria to avoid collinearity of VAs (Sec  IV-B2). Next, we propose a non-uniform continuous-time state estimation method, using the adaptive knot span adjustment strategy to tightly fuse the high-frequency measurements from UWB ranging, IMU, and odometer (Sec. IV-B3). The backend conducts a UWB-Inertial-Odometer factor graph with an adaptive sliding window to perform joint optimization to obtain optimal estimations (Sec. IV-C).
For the UWB/IMU/Odometer (UIO) integrated system, we use U{U} to denote the UWB frame, $I$ for the IMU frame, $O$ for the wheel odometer frame, $W$ for the world coordinate frame.
IV-A Preprocessing
IV-A1 UWB Ranging Outlier Rejection
$\left|{\left\|\hat{\textbf{x}}_{t_{k}}-\textbf{p}_{n}\right\|}_{2}-r_{t_{k+1}}\right|<\gamma\hat{v}_{t_{k}}\left(t_{k+1}-t_{k}\right),$ (1)
where $\hat{\textbf{x}}_{t_{k}}$ and $\hat{v}_{t_{k}}$ are the estimated robot’s position and velocity derived from the IMU/odometer fusion model, respectively. $\textbf{p}_{n}$ is the position of the UWB anchor $n$. $\gamma$ is the outlier rejection gating threshold for the individual ranging measurement.
IV-A2 IMU/Odometer Fusion Model
@@
IMU and odometer offer fast response speeds and are environment-independent. For ground robot navigation, it appears to be straightforward to fuse IMU data with wheel odometer measurements to provide motion priors over short time periods [Bai_Lai_Lyu_Cen_Wang_Sun_2022].
The raw IMU measurements include the angular velocity $\omega_{t_{k}}$ and local linear acceleration $a_{t_{k}}$:
$^{I}\omega_{t_{k}}=\tilde{{}^{I}\omega_{t_{k}}}+\varepsilon_{g}+n_{g}$ (2)
$^{I}a_{t_{k}}=\tilde{{}^{I}a_{t_{k}}}+\varepsilon_{a}+\prescript{I}{W}{\textbf{R}}g^{W}+n_{a},$ (3)
@@
where $\tilde{\omega}_{t_{k}}$ and $\tilde{a}_{t_{k}}$ are true values of angular velocity and linear acceleration at time $t_{k}$, respectively. $\varepsilon_{g}$ and $\varepsilon_{a}$ denote the accelerometer’s bias and the gyroscope’s bias in the IMU body frame. $\prescript{I}{W}{\textbf{R}}$ denotes the rotation matrix from frame $\left\{W\right\}$ to $\left\{I\right\}$, $g^{W}$ is the gravity vector in the world frame, $n_{g}$ and $n_{a}$ are zero-mean Gaussian noises.
We then integrate the wheel odometer measurements between IMU times $t_{k-1}$, $t_{k}$] and 2D relative motion as follows:
$z_{o}=:\begin{bmatrix}{}^{o}\theta\\ {}^{o}\textbf{d}\end{bmatrix}=\begin{bmatrix}-\int\limits_{t\in[t_{k-1},t_{k}]}{{}^{o}\omega_{t_{k}}dt}\\ \int\limits_{t\in[t_{k-1},t_{k}]}{{}^{o}v_{t_{k}}\text{cos}^{o}\theta dt}\\ -\int\limits_{t\in[t_{k-1},t_{k}]}{{}^{o}v_{t_{k}}\text{sin}^{o}\theta dt}\end{bmatrix}$, (4)
@@
$\tilde{z}_{o}=:\begin{bmatrix}{}_{I}^{o}\theta\\ {}_{I}^{o}\textbf{d}\par\end{bmatrix}=\begin{bmatrix}\textit{Log}\left(\textbf{e}_{3}^{\text{T}}{{}^{o}R_{t_{k}}}\right)\\ {{}^{o}R_{t_{k}}}(_{I}^{o}R{{}^{I}\textbf{d}})\end{bmatrix}$. (5)
$\textit{Log}(\cdot)$ is the matrix logarithm function, $\textbf{e}_{3}$ is the unit bias vector. The attitude update is formulated as
@@
$^{o}R_{t_{k}}$} represents the rotation matrices transformed from $^{o}\theta_{t_{k}}$. We construct the observation measurement vector composed of the poses of the IMU and the odometer as
@@
where $\eta(k)\sim(0,\sigma^{2})$ means the measurement noise with zero mean and a standard Gaussian distribution. The Jacobian matrix $H(k)$ can be indicated by
@@
The innovation $\tilde{Z}(k)$ can be calculated as
@@
Here, $M$ is the adaptive size of the observation statistics, which controls the trade-off between estimation accuracy and computational complexity. $\mu$ is the convergence rate. $\alpha$ is a constant associated with the frequency characteristics of the IMU measurement noise. $\xi$ is the maximum allowable size of $M$. $\lambda_{\text{min}}$ and $\lambda_{\text{max}}$ are lower and upper thresholds for the acceptable range of $e(k)$. There is the rejection domain of $E(k)$:
$\left\{\begin{matrix}E(k)>thr&\text{Reject}\\ E(k)\leq thr&\text{Accept}\end{matrix}\right.$. (15)
@@
When $E(k)$ exceeds the threshold thrthr, the abnormal fault in IMU measurements is detected by the chi-square test statistic, and we reject this IMU measurement and do not perform the EKF update.
IV-B Frontend
IV-B1 Virtual Anchor (VA) Generation
Since directly fusing range measurements can lead to unobservability when only one or two anchors are available [10268655], we utilize virtual anchors (VAs) to construct a fully observable UWB-based localization system. Inspired by [9119679, 9520065], we propose a VA generation method based on multiple hypotheses. This method introduces location reference points that are virtually placed in the UWB network, providing additional constraints for the robot’s global localization, as shown in Fig. 2. The VA is established by a location estimator from the IMU/odometer fusion model. In this article, considering a mobile robot equipped with a UWB tag that receives a signal from a UWB anchor, the corresponding VAs can be generated by range measurements and the estimated position using the IMU/odometer fusion model. The range residual is established as
$\textbf{e}_{j}^{r}=r_{j}-\left\|\textbf{x}_{j}-\textbf{p}_{n}\right\|_{2},$ (16)
Figure 2: Illustrations of the VA generation process.
Since $\textbf{q}^{\omega}$ and $\triangle\textbf{d}_{ij}$ can be derived from Eq. (4), the robot’s pose and the derived relative position can be expressed as the VA’s position. Eq. (16) can be further reformulated as follows:
where $\hat{\textbf{p}}_{n}^{j}$ is regarded as the position of the VA. Based on this, the VA is generated through the ranging-gathering process that involves a set of robot waypoints. However, the random waypoints are often ill-posed in practice, resulting in suboptimal solutions. Selecting an appropriate information metric to detect potential sensor measurement degeneration and determining the optimal waypoint placement are critical for enhancing localization accuracy. The Fisher information Matrix (FIM) [6882246] is considered to evaluate the performance of VA’s position estimation, and it is defined as:
$\textbf{F}=\frac{1}{\sigma_{r}^{2}}\textbf{J}^{\text{T}}\textbf{J},$ (19)
where $\sigma_{r}$ is the standard deviation of the ranging measurements. J represents the Jacobian matrix of the ranging measurement model with respect to the anchor’s position $\hat{\textbf{p}}_{n}^{j}$. The determinant of the FIM $det(\textbf{F})$ contains all the ranging measurements received during time $i$ to $j$. We can set a threshold $\tau_{F}$ and generate a VA when
$det(\textbf{F})>\tau_{F}.$ (21)
By maximizing $det(\textbf{F})$, the lower bound of the VA’s position estimation error can be directly determined, and the optimal VA placement can be decided. The VA generation process continues until three non-collinear VAs are obtained.
IV-B2 Multiple Hypotheses-based Criteria for VAs
We present multiple hypotheses-based criteria to avoid collinearity of generated VAs. A set of VAs is derived from the IMU/odometer fusion model. If all the VAs’ 2D positions constitute a straight line, it becomes difficult to obtain a unique solution for the target’s position without additional constraints. To mitigate this problem, we use the determinant to ensure that the generated VAs are non-collinear. Assuming the locations of three VAs are $\hat{\textbf{p}}_{n}^{1}=(\hat{x}_{n}^{1},\hat{y}_{n}^{1}),\hat{\textbf{p}}_{n}^{2}=(\hat{x}_{n}^{2},\hat{y}_{n}^{2}),\hat{\textbf{p}}_{n}^{3}=(\hat{x}_{n}^{3},\hat{y}_{n}^{3})$. The determinant $D$ is defined as:
$D=\begin{vmatrix}\hat{x}_{n}^{1}&\hat{y}_{n}^{1}&1\\ \hat{x}_{n}^{2}&\hat{y}_{n}^{2}&1\\ \hat{x}_{n}^{3}&\hat{y}_{n}^{3}&1\end{vmatrix}$. (22)
If $D=0$, the three VAs are collinear. We use a sample circle with multiple VAs’ position hypotheses to avoid collinearity. Thus, assuming total $h$ different hypotheses for position $\hat{\textbf{p}}_{n}^{j}$, the VA’s position will be written in the following form:
$\textbf{b}_{n}^{j}=\textbf{H}_{j}^{k}\hat{\textbf{p}}_{n}^{j},$ (23)
$\textbf{H}_{j}^{k}=\begin{bmatrix}r_{j}&cos(2\pi(k-1)/h)&0\\ 0&r_{j}sin(2\pi(k-1))/h)&0\\ 0&0&1\end{bmatrix},k=1,...,h.$ (24)
The resultant VAs’ positions are obtained by minimizing the following expression:
$\text{arg min}\sum_{j=1}^{h}(r_{j}-\left\|\textbf{x}_{i}-\textbf{b}_{n}^{j}\right\|).$ (25)
IV-B3 Adaptive Knot Span Adjustment Strategy
@@
In this article, we use the split non-uniform B-spline representation to separately represent the translation part $\textbf{cp}(t)$ in $\mathbb{R}^{3}$ and the rotation part $\textbf{R}(t)$ in the Lie group $SO(3)$ with two splines. The sampled translation vector in trajectory of the order $k$ B-spline at time $t\in[t_{k},t_{k+1}]$, can be denoted as:
$\textbf{cp}(t)=\textbf{cp}_{i}+\sum_{j=1}^{k}\lambda_{j}(t)\cdot\textbf{d}_{j}^{i},$ (26)
@@
where $\textbf{cp}_{i}$ denotes the $i$-th control point, $\textbf{cp}_{i}\in\mathbb{R}^{3}$, $t\in T=\left\{t_{0},t_{1},...,t_{n}\right\}$ is the arbitrarily independent knot time of B-splines. The coefficients ${\lambda}_{j}(t)$ represents the $j$th column of $\boldsymbol{\lambda}(t)$. $\tilde{\boldsymbol{\Lambda}}^{(k+1)}(i)$ is a non-constant cumulative basis matrix which is defined by its order $k$. $\textbf{u}(t)=\begin{bmatrix}1&u(t)&...&u^{k}(t)\end{bmatrix}^{\text{T}}$ is the normalized time vector:
@@
$\textbf{d}_{j}^{i}=\textbf{cp}_{i+j}-\textbf{cp}_{i+j-1}.$ (29)
@@
To balance the overall accuracy and convergence speed, we exploit a cumulative cubic B-spline function (k=3k{=}3) in this work, and the order of the spline is set to $4$.
For uniform-spline-based methods, control points are temporally uniformly distributed, and knot spans between them are equal to construct the B-spline. However, this uniform placement does not guarantee optimal trajectory interpolation performance, especially when encoding time-varying states. We argue that it is reasonable to reduce redundant control points for steady movement. In addition, when the robot’s motion has significant velocity fluctuations over short time intervals, a finer granularity of knots is required to capture rapid changes, which requires inserting additional control points. To solve this problem, we propose an adaptive knot span adjustment strategy that dynamically adjusts the distribution of control points to flexibly accommodate varying motion speeds, as shown in Fig. 3.
((a))
((b))
Figure 3: Illustrations of the proposed adaptive knot span adjustment strategy. The top shows the robot’s velocity, and the bottom shows that the distribution of control points is adaptive to speed variations.
Namely, the current pose is registered as a Key Pose (KP), if the tracked trajectory segment has a local translational velocity or rotational velocity change above a certain threshold. Specifically, we obtain the translational velocity and rotational velocity respectively in a sliding window:
$\Delta v(T_{i-1},T_{i})$ $=\frac{1}{n(T_{i-1},T_{i})}\sum_{k\in n(T_{i-1},T_{i})}$ (31)
@@
$\Delta\omega(T_{i-1},T_{i})$ $=\frac{1}{n(T_{i-1},T_{i})}\sum_{k\in n(T_{i-1},T_{i})}$
$\Big(\omega_{k}$ $-\frac{1}{n(T_{i-1},T_{i})}\sum_{k\in n(T_{i-1},T_{i})}\omega_{k}\Big)^{2},$
where $n(T_{i-1},T_{i})$ represents the number of samples within a sliding window. We associate each KP in $T_{i-1},T_{i}]$ with the number of control points $n_{cp}$
@@
where $\lambda_{1}$, $\lambda_{2}$, $\lambda_{3}$, $\lambda_{4}$ define the translational-velocity-change bins, and $\theta_{1}$, $\theta_{2}$, $\theta_{3}$, $\theta_{4}$ define the rotational-velocity-change bins. We know that the knot span is actually influenced by the number of control points at a specific time interval. Each of the knot spans is independent of the others, and the knot spans are divided by control points:
$\bigtriangleup t_{i}=\frac{1}{n_{cp}}(T_{i}-T_{i-1}).$ (33)
@@
Then, the trajectory segment over the sliding window $T_{i-1},T_{i}]$ consists of $n_{cp}$ piecewise functions. We can query the pose with an arbitrary timestamp $t_{i}$ in this segment. The control points of the active trajectory segment are further optimized and updated within the sliding window by a factor graph optimization. Meanwhile, control points outside the active segment remain static in the optimization but still retain marginalization priors for the next sliding window.
IV-C Backend
IV-C1 Virtual Anchor Ranging Factor
Through constructing three non-collinear VAs, a global position $\textbf{x}_{t}$ of the robot at time $t$ can be obtained. Then, the residual of the virtual anchor ranging factor can be written as:
$\textbf{e}_{t}^{r}=\sum_{m\in M}r_{t}^{m}-\left\|\textbf{x}_{t}-\hat{\textbf{p}}_{n}^{m}\right\|_{2},$ (34)
where $M$ is the set of location information $\hat{\textbf{p}}_{n}^{m}$ of the VAs detected at time $t$.
IV-C2 IMU Factor